Determine the truth value of the following statement.

(∃n ∈Z)(n2= 3)



1
Expert's answer
2022-02-07T17:25:21-0500

Solution:

"(∃n ∈Z)(n^2= 3)
\\\Rightarrow n^2=3
\\\Rightarrow n=\pm \sqrt 3 \notin Z"

But given is "n \in Z".

Thus, it is false.

Truth value is F.